OpenStudy (anonymous):

Solve by using the quadratic formula to solve x2 + 6x = -3.

OpenStudy (arabpride):

is it x^2

OpenStudy (anonymous):

oops! sorry! Yeah I meant x²

OpenStudy (arabpride):

lol yuperxzzz close enough XD

OpenStudy (anonymous):

hahaha :)

OpenStudy (johnweldon1993):

Add 3 to both sides and set = to 0 \[\large x^2 + 6x + 3 = 0\] Coefficients are a = 1 b= 6 c = 3 \[\large \frac{-b \pm \sqrt{b^2 - 4ac)}}{2a}\] Plug in your numbers and solve :)

OpenStudy (arabpride):

BTW: WELCOME TO OPENSTUDY!!!!! *^.^*

OpenStudy (anonymous):

aww thank you! Its amazing so far! :)

OpenStudy (anonymous):

Thank you John!

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (arabpride):

well maybe uve only been on for like an hour XD

OpenStudy (anonymous):

hahaha

OpenStudy (arabpride):

But if u need anyhelp u can jus message me or tag me Tag people by pressing "@' then tyoe in the user name of the person u need help from, after u fan them *^.^*

OpenStudy (anonymous):

okay :) Thanks so much!!

OpenStudy (arabpride):

UR VERY WELCOME!!! ANYTIME!!! *^.^*